Mr B. deposits at the end of each year 3,000 euros with an annual interest rate of 2% and this is continued for 6 years. The capital formed at the end of year 6 is deposited in the same bank with a compound interest and the annual interest rate is 1.5%. Find the amount that Mr B. will receive three years after the last deposit?
Answer #1
Step-1:Future value of annual deposit for 6 years
Future value = Annual deposit * Future value of annuity of 1
= €          3,000 * 6.308121
= €       18,924
Working:
Future value of annuity of 1 = (((1+i)^n)-1)/i Where,
= (((1+0.02)^6)-1)/0.02 i = 2%
= 6.30812096 n = 6
Step-2:Future value of investment 3 years after last deposit
Future value = P*(1+i)^n Where,
= 18924.36*(1+0.015)^3 P = € 18,924.36
= €       19,789 i = 1.50%
n = 3